# Resource-Adaptive Successive Doubling for Hyperparameter Optimization with Large Datasets on High-Performance Computing Systems

### v1

@@ Abstract
journal: Future Generation Computer Systems
@@ 1 Introduction
The developed method is suitable for problems that involve large scientific datasets, where due to long training times, even with HPC resources it is not feasible to train more than the initially sampled hyperparameter configurations and users are interested in getting the best possible, fully-trained model in the shortest amount of time. Therefore, this study performs an extensive evaluation of RASDA on different datasets from the CV, CFD, and AM domains, which are up to 8.3  Terrabyte (TB) in size, to prove its capability to deal with large datasets. These datasets are used to tune the hyperparameters of different types of NNs, namely a Convolutional Neural Network (CNN), an autoencoder and a transformer. RASDA is also benchmarked against the current state-of-the-art successive halving HPO method ASHA. The new RASDA code is openly available on GitHub [^1] RASDA source: https://github.com/olympiquemarcel/rasda for the community.
@@ 3 Resource-Adaptive Successive Doubling Algorithm
This section presents details on RASDA in Sec. 3.1 and provides an explanation on how issues with large batch size training, cf. Sec. 2.4, are addressed in Sec. 3.2.
@@ 3.2 Large Batch Training
McCandlish et al. [39] empirically studied large batch training for various models: they introduce the Gradient Noise Scale (GNS) metric, which serves as a noise-to-signal measure of the training progress. In theory, if the true gradient $G_{true}$ from performing full-batch Gradient Descent without the stochastic component would be available, it would be possible to compute the a simple version of the GNS by
@@
Another crucial point is the correct scaling of the learning rate with the batch size. In the evaluation in Sec. 5, the learning rate is scaled linearly with the number of workers, i.e., up to a factor of 8×8$\times$8 ×, when using SGD [28]. Furthermore, it follows a square-root scaling rule when using Adaptive Moment Estimation (ADAM[37]. In the case of re-scaling, the learning rate is not immediately scaled to a larger value. Instead, there is a warm-up over one or two epochs. This re-scaling parameter is included as a hyperparameter in the search space, see Tab. 1. Thereby, the HPO run automatically optimizes towards learning stability.
@@ 4 Application Cases
To assess the proposed RASDA scheduler, its performance is evaluated across a range of different tasks from the CV, CFD, and AM domain. These cases feature various models with different hyperparameters to optimize as well as training datasets of different sizes. The application domains and the set-up of these tasks is described in the following.
@@
Learning rateCV, CFD, ${}^{\text{CV, CFD, AM}}$ float log[1e-5, 1]
Weight decayCV, CFD, ${}^{\text{CV, CFD, AM}}$ float log${(0,1e-1]}$
Initial warm-upCV, CFD, ${}^{\text{CV, CFD, AM}}$ int [1,2,3,4,5]
OptimizerCV, CFD, ${}^{\text{CV, CFD, AM}}$ cat [”sgd”, ”adam”]
Layer initializationCV, ${}^{\text{CV, CFD}}$ cat [”kaiming” [[15]],
@@
Convolution kernel sizeCV, ${}^{\text{CV, CFD}}$ int [5,7,9]
Re-scaling warm-upCFD, ${}^{\text{CFD, AM}}$ int [1,2]
@@ 4.1 Computer Vision
All models are trained for $\min\_t=5$ to $\max\_t$ $=40$ epochs and a reduction and scaling factor $\mathtt{sf}=\mathtt{rf}=2$ is chosen for the schedulers. Following Eq. (3), this results in rung milestones at epochs $5,10,20,$ and $40$.
@@ 4.2 Additive Manufacturing
The AM dataset is taken from the RAISE-LPBF benchmarking dataset [9], which includes a selection of high-speed video recordings at 20,000 frames per second of a laser powder bed fusion processes for stainless steel. The laser power and speed parameters are systematically varied. The goal is to reconstruct the power and speed of the laser from this video input. By comparing the predicted laser parameters with the pre-set parameters of the machine producing the laser, anomalies in the printing process can be detected faster, leading to more efficient quality control. The base ML model used for this task is a SwinTransformer [47], with the HPO search space consisting of multiple, Transformer-specific architectural and optimizer-related parameters, such as the number of attention heads, see Tab. 1. The model is trained on the C027 cylinder with a 80/20 split for training and validation and is approximately 60 GB in size. It is evaluated on the C028 cylinder for testing purposes. The Mean-Squared Error (MSE) between predicted and actual laser power and speed is computed to assess the accuracy of the SwinTransformer. All models are trained for $\min\_t=5$ to $\max\_t$ $=20$ epochs and a reduction and scaling factor $\mathtt{sf}=\mathtt{rf}=2$ is chosen for the schedulers. Following Eq. (3), this results in rung milestones at epochs $5,10,$ and $20$.
@@ 4.3 Computational Fluid Dynamics
The CFD dataset contains actuated turbulent boundary layer flow data, generated from a simulation [4]. The CFD dataset is stored in HDF5 file format and comprises several widths. In this study, widths of $1{,}000,1{,}200,$ and $1{,}600$ are used as training dataset (approximately $4.8$ TB in size). Width of $1{,}800$ and $3{,}000$ are used as validation and test datasets (approximately $3.5$ TB in size) to assess extrapolation performance. Altogether, the dataset is approximately $8.3$ TB in size.
@@ 5 Results
This section presents the experimental results of running the proposed algorithm on two supercomputer systems, which are introduced in Sec. 5.1. Section 5.2 focuses on the scaling performance of the RASDA algorithm on up to 1,024 GPUs, while Sec. 5.3 compares the RASDA against the plain ASHA scheduler without any resource adaptation.
@@ 5.2 Scaling Performance
The first weak scaling experiment considers a smaller scale of $8$ to $64$ GPUs. The runtime and accuracy of the RASDA algorithm is compared to the plain ASHA algorithm for training a ResNet50 on the ImageNet dataset for 20 epochs, see Fig. 3. It can be seen that on all scales (from 8 to 64 GPUs), the RASDA algorithm achieves consistently lower runtimes up to a factor of $1.45$ faster than its ASHA counterpart while matching the final test set accuracy in almost all cases.
The second scaling experiment considers a large scale of $128$ to $1{,}024$ GPUs, see Fig. 4. The weak scalability of the RASDA algorithm is evaluated by training a ResNet for six epochs. The results show that the algorithm maintains a high parallel efficiency of $>0.84$ on up to $1{,}024$ GPUs.
@@
Figure 4: Weak scalability of the RASDA algorithm on up to $1{,}024$ GPUs on the JUWELS BOOSTER system, including ideal scalability for comparison.
@@ 5.3 Speed-Ups and Accuracy
Train Accuracy $\uparrow$ 0.6976 0.7310 1.05×1.05$\times$1.05 ×
Val Accuracy $\uparrow$ 0.6728 0.6813 1.01×1.01$\times$1.01 ×
Test Accuracy $\uparrow$ 0.6688 0.6766 1.01×1.01$\times$1.01 ×
Runtime (in seconds) $\downarrow$ 31637 18502 1.71×1.71$\times$1.71 ×
@@
Val MSE $\downarrow$ 0.0455 0.0404 1.12×1.12$\times$1.12 ×
Test MSE $\downarrow$ 0.0554 0.0516 1.07×1.07$\times$1.07 ×
Runtime (in seconds) $\downarrow$ 5784 3803 1.52×1.52$\times$1.52 ×
@@
[…]
Test Relative Error $\downarrow$ 0.0185 0.0115 1.61×1.61$\times$1.61 ×
Runtime (in seconds) $\downarrow$ 19487 10242 1.90×1.90$\times$1.90 ×
@@ 5.4 Performance at 1,024 GPUs Scale
[…]
@@
[…]
Test Rel. Error 0.0016 7.17×7.17$\times$7.17 ×
@@ 6 Summary and Outlook
In addition, this study represents the first application of systematic HPO to a scientific dataset at the TB scale. A comparison of the application of RASDA on 128 and 1024 GPUs revealed a significant improvement in model performance. Specifically, the larger-scale application identifies a model that is significantly superior in solution quality. These results demonstrate the scalability and efficiency of the RASDA method, thus paving the way for the application of HPO methods on current and future Exascale supercomputers. For future work, the optimal timing for scaling the batch size (through the addition of more GPUs to the data-parallel training loop) should be investigated more thoroughly. Furthermore, the impact of scaling the batch size on various hyperparameters beyond the learning rate (such as the weight decay values) warrants a deeper exploration.

### v2

@@ Abstract
journal: Future Generation Computer Systemsfootnotetext: © 2025. This manuscript version is made available under the $C$-BY-NC-ND 4.0 license https://creativecommons.org/licenses/by-nc-nd/4.0/.
@@ 1 Introduction
The developed method is suitable for problems that involve large scientific datasets, where due to long training times, even with HPC resources it is not feasible to train more than the initially sampled hyperparameter configurations and users are interested in getting the best possible, fully-trained model in the shortest amount of time. Therefore, this study performs an extensive evaluation of RASDA on different datasets from the CV, CFD, and AM domains, which are up to 8.3  Terrabyte (TB) in size, to prove its capability to deal with large datasets. These datasets are used to tune the hyperparameters of different types of NNs, namely a Convolutional Neural Network (CNN), an autoencoder and a transformer. RASDA is also benchmarked against the current state-of-the-art successive halving HPO method ASHA. The new RASDA code is openly available on GitHub [^1] RASDA source: https://github.com/olympiquemarcel/rasda for the community, see Tab. 8 for an overview of the repository.
In summary, the key contributions of RASDA are:
  • 1.
    Combination of successive halving in space with successive doubling in time, allocating more GPUs to more promising trials.
  • 2.
    Reduces the runtime of more promising hyperparameter trials by leveraging a higher degree of parallelism in data-parallel training
  • 3.
    Leverages the inherent features of HPC systems, fast computation for the training, and fast communication for exchange of gradients during distributed training
  • 4.
    Outperforms the plain ASHA method in runtime and model performance across different domains and on datasets up to 8.3 TB.
@@ 3 Resource-Adaptive Successive Doubling Algorithm
This section presents details on RASDA in Sec. 3.1 and provides an explanation on how issues with large batch size training, cf. Sec. 2.4, are addressed in Sec. 3.2. The performance optimizations are presented in Sec. 3.3, while in Sec. 3.4 and Sec. 3.5 the compatibility of RASDA with other tools and its dependency on the hardware setup are summarized.
@@ 3.2 Large Batch Training
McCandlish et al. [41] empirically studied large batch training for various models: they introduce the Gradient Noise Scale (GNS) metric, which serves as a noise-to-signal measure of the training progress. In theory, if the true gradient $G_{true}$ from performing full-batch Gradient Descent without the stochastic component would be available, it would be possible to compute a simple version of the GNS by
@@
Another crucial point is the correct scaling of the learning rate with the batch size. In the evaluation in Sec. 5, the learning rate is scaled linearly with the number of workers, i.e., up to a factor of $8\times$, when using SGD [29]. Furthermore, it follows a square-root scaling rule when using Adaptive Moment Estimation (ADAM[38]. In the case of re-scaling, the learning rate is not immediately scaled to a larger value. Instead, there is a warm-up over one or two epochs. This re-scaling parameter is included as a hyperparameter in the search space, see Tab. 1. Thereby, the HPO run automatically optimizes towards learning stability.
@@ 3.3 Performance Optimization
3.4 Compatibility with Existing HPO Tools
As RASDA functions as a pure scheduling tool, it can be integrated with various HPO and AutoML frameworks. Since it is already incorporated into the Ray Tune framework via the scheduler interface, it can be seamlessly used within workflows that leverage the scheduling infrastructure of Ray Tune. This includes BO tools such as BOHB [12], Optuna [3], and BayesOpt [39], as well as evolutionary optimization frameworks like HEBO [10]. In such cases, the Bayesian or evolutionary algorithm proposes new hyperparameter candidates to sample from the search space, while RASDAs time-wise successive halving determines which trials to terminate at different points in time. Concurrently, its space-wise successive doubling allocates additional GPU resources to the trials selected to continue.
3.5 Dependency on Hardware Setup
As the main idea of RASDA is to perform successive doubling in the spatial domain, the scheduler operates most efficiently on HPC systems where the number of GPUs per node, the total number of allocated nodes, and the number of concurrent trials follow a power-of-two configuration. This setup ensures that GPUs resources can be reassigned seamlessly across rung milestones, as illustrated in Fig. 2. However, such configurations may not always be available in practice. In scenarios where the number of GPUs is not a power of two or where the number of trials exceeds the number of available GPUs, RASDA relies on the underlying Ray Tune framework to manage resource scheduling and queuing. When more trials are submitted than there are GPUs available, not all trials can be launched simultaneously. In this case, RASDA proceeds with its successive doubling routine for the first batch of trials that are scheduled. Remaining trials are queued and executed as resources become available. At the milestones, trials selected to continue are paused if their requested GPU allocation cannot be satisfied immediately, and they are resumed once sufficient resources are freed by completed or terminated trials. Although this queuing may introduce some latency, the promoted trials benefit from increased parallelism once resumed, resulting in significantly faster training. The benefit becomes especially pronounced in higher rungs, where larger resource allocations substantially reduce the training time per epoch. As a result, RASDA is still expected to deliver well-performing hyperparameter candidates faster than plain ASHA. Similarly, in cases where the total number of GPUs is not a power of two, the successive doubling scheme can still be applied, although some adaptation is required. Specifically, the values of $\mathtt{sf}$, $\mathtt{rf}$, $\min\_t$, and $\max\_t$ should be chosen to ensure that the maximum number of GPUs allocated per trial in the final rung does not exceed the available resources. It is generally advisable to avoid resource fragmentation across nodes, as splitting the GPUs on a node between multiple trials may reduce the efficiency of data-parallel training.
@@ 4 Application Cases
To assess the proposed RASDA scheduler, its performance is evaluated across a range of different tasks from the CV, CFD, and AM domain on seperate training, validation and test dataset splits to avoid overfitting. The different cases feature various models with different hyperparameters to optimize as well as training datasets of different sizes. The application domains and the set-up of these tasks is described in the following.
@@
[…]
@@
[…]
@@ 4.1 Computer Vision
All models are trained for $\min\_t=5$ to $\max\_t$ =40=40 epochs and a reduction and scaling factor $\mathtt{sf}=\mathtt{rf}=2$ is chosen for the schedulers. Following Eq. (3), this results in rung milestones at epochs $5,10,20,$ and $40$.
@@ 4.2 Additive Manufacturing
The AM dataset is taken from the RAISE-LPBF benchmarking dataset [9], which includes a selection of high-speed video recordings at 20,000 frames per second of a laser powder bed fusion processes for stainless steel. The laser power and speed parameters are systematically varied. The goal is to reconstruct the power and speed of the laser from this video input. By comparing the predicted laser parameters with the pre-set parameters of the machine producing the laser, anomalies in the printing process can be detected faster, leading to more efficient quality control. The base ML model used for this task is a SwinTransformer [49], with the HPO search space consisting of multiple, Transformer-specific architectural and optimizer-related parameters, such as the number of attention heads, see Tab. 1. The model is trained on the C027 cylinder with a 80/20 split for training and validation and is approximately 60 GB in size. It is evaluated on the C028 cylinder for testing purposes. The Mean-Squared Error (MSE) between predicted and actual laser power and speed is computed to assess the accuracy of the SwinTransformer. All models are trained for $\min\_t=5$ to $\max\_t$ =20=20 epochs and a reduction and scaling factor $\mathtt{sf}=\mathtt{rf}=2$ is chosen for the schedulers. Following Eq. (3), this results in rung milestones at epochs $5,10,$ and $20$.
@@ 4.3 Computational Fluid Dynamics
The CFD dataset contains actuated turbulent boundary layer flow data, generated from a simulation [4]. The CFD dataset is stored in HDF5 file format and comprises several widths. In this study, widths of 1,000,1,200,1{,}000,1{,}200, and 1,6001{,}600 are used as training dataset (approximately $4.8$ TB in size). Width of 1,8001{,}800 and 3,0003{,}000 are used as validation and test datasets (approximately $3.5$ TB in size) to assess extrapolation performance. Altogether, the dataset is approximately $8.3$ TB in size.
@@
All experiments use reduction and scaling factors $\mathtt{sf}=\mathtt{rf}=2$, as this provides a suitable trade-off between terminating unpromising trials and scaling up GPU resources. Choosing a scaling factor that is too large can lead to learning instabilities at the rung milestones, due to the abrupt increase in the number of GPUs allocated. Similarly, selecting a reduction factor that is too large may result in prematurely terminating trials that could have shown strong performance in later training stages. The value $\min_{t}=5$ is selected to ensure sufficient initial training and to avoid interference with the learning rate warm-up phase, which can last up to five epochs (see Tab. 1). Terminating trials during this period could result in inaccurate early stopping decisions. The values $\max_{t}=20$ and $\max_{t}=40$ are chosen to provide adequate training time for convergence, while also considering computational resource constraints.
@@ 5 Results
This section presents the experimental results of running the proposed algorithm on two supercomputer systems, which are introduced in Sec. 5.1. Section 5.2 focuses on the scaling performance of the RASDA algorithm on up to 1,024 GPUs, while Sec. 5.3 compares the RASDA against the plain ASHA scheduler without any resource adaptation. Section 5.4 reports the performance of RASDA at the large scale, and in Sec. 5.5 different ablation experiments are presented.
@@ 5.2 Scaling Performance
The first weak scaling experiment considers a smaller scale of $8$ to $64$ GPUs. The runtime and accuracy of the RASDA algorithm is compared to the plain ASHA algorithm for training a ResNet50 on the ImageNet dataset for 20 epochs, see Fig. 3. It can be seen that on all scales (from 8 to 64 GPUs), the RASDA algorithm achieves consistently lower runtimes up to a factor of $1.45$ faster than its ASHA counterpart while matching the final test set accuracy in almost all cases. Reduced test set accuracy is observed only at $8$ and, to a lesser extent, $16$ GPUs, as in these small settings the number of hyperparameter configurations evaluated is low and RASDA cannot yet fully benefit from large batch size training (see Sec. 3.2).
The second scaling experiment considers a large scale of $128$ to 1,0241{,}024 GPUs, see Fig. 4. The weak scalability of the RASDA algorithm is evaluated by training a ResNet for six epochs. The results show that the algorithm maintains a high parallel efficiency of >0.84>0.84 on up to 1,0241{,}024 GPUs.
@@
The better scaling performance of RASDA compared to ASHA can also be observed when examining specific HPC metrics. RASDA consistently achieves higher GPU utilization, as it avoids idling GPUs, see Fig. 2. RASDA achieves approximately $\approx 80\%$ GPU utilization, compared to around $\approx 54\%$ for the ASHA case. Since more GPUs are actively used, this also leads to higher I/O demand compared to ASHA, where idle GPUs consume less data. However, because RASDA achieves significantly shorter runtimes, it is expected to be more energy-efficient overall.
@@
Figure 4: Weak scalability of the RASDA algorithm on up to 1,0241{,}024 GPUs on the JUWELS BOOSTER system, including ideal scalability for comparison.
@@ 5.3 Speed-Ups and Accuracy
Train Accuracy $\uparrow$ 0.6976 0.7310 $1.05\times$
Val Accuracy $\uparrow$ 0.6728 0.6813 $1.01\times$
Test Accuracy $\uparrow$ 0.6688 0.6766 $1.01\times$
Runtime (in seconds) $\downarrow$ 31637 18502 $1.71\times$
@@
Val MSE $\downarrow$ 0.0455 0.0404 $1.12\times$
Test MSE $\downarrow$ 0.0554 0.0516 $1.07\times$
Runtime (in seconds) $\downarrow$ 5784 3803 $1.52\times$
@@
Val MSE $\downarrow$ $5.28\text{\times}{10}^{-06}$ $2.81\text{\times}{10}^{-06}$ $1.88\times$
Test MSE $\downarrow$ $4.42\text{\times}{10}^{-06}$ $2.40\text{\times}{10}^{-06}$ $1.84\times$
Test Relative Error $\downarrow$ 0.0185 0.0115 $1.61\times$
Runtime (in seconds) $\downarrow$ 19487 10242 $1.90\times$
@@ 5.4 Performance at 1,024 GPUs Scale
While the superiority of RASDA over plain ASHA has been confirmed in the previous experiments using 64 and 128 GPUs, a final RASDA experiment on a 1,024 GPU scale is conducted on the JUWELS BOOSTER system. Again, using the CFD application case, the number of configurations to be evaluated is increased to 64, with each trial starting with 16 GPUs. The models are trained for $\min\_t=5$ and $\max\_t=20$ epochs. The HPO run took three hours and resulted in an improved model with a validation MSE of $\approx$3.63\text{\times}{10}^{-07}$\$, a test MSE of $\approx$4.88\text{\times}{10}^{-08}$\$ and a relative test error of $\approx 0.0016$. Depending on the metric, this is a $7$ to $49$ times increase in solution quality, compared to the results of the HPO run on 128 GPUs (see Tab. 5), which highlights the potential of large-scale HPO for scientific ML.
@@
Val MSE $3.63\text{\times}{10}^{-07}$ $7.74\times$
Test MSE $4.88\text{\times}{10}^{-08}$ $49.22\times$
Test Rel. Error 0.0016 $7.17\times$
5.5 Ablation Studies
To evaluate the impact of different parameters on the performance of the RASDA method, two ablation studies are conducted.
5.5.1 Impact of Reduction and Scaling Factors
All prior experiments use a scaling and reduction factor of $\mathtt{sf}=\mathtt{rf}=2$. In this ablation study on the CV application case, these factors are increased to $\mathtt{sf}=\mathtt{rf}=4$. Since the number of GPU accelerators per node on HPC systems typically follows a power-of-two configuration, allocating partial nodes may lead to performance degradation. To ensure sufficient training time between decision points, the same values of $\min\_t=5$ and $\max\_t=20$ epochs are retained. According to Eq. (3), this results in two rung milestones at epochs $5$ and $20$. As shown in Tab. 6, RASDA continues to outperform ASHA under these settings, achieving higher accuracy and a runtime speed-up of $\approx 1.6\times$. However, this speed-up is smaller than the improvement observed with $\mathtt{sf}=\mathtt{rf}=2$ (see Tab. 2). While a larger scaling factor enables more aggressive allocation of GPU resources to promising trials, it also increases the time between rung milestones due to the geometric progression (see Eq. 3). This observation indicates that $\mathtt{rf}=2$ is a suitable choice for the RASDA scheduler.
Table 6: Comparison of ASHA and RASDA with $\mathtt{sf}=\mathtt{rf}=4$ for the CV application case, trained for 20 epochs on 64 GPUs on the JURECA-DC-GPU system. Results are averaged over three random seeds. Better results ($\uparrow$ or $\downarrow$ depending on the metric) are underlined.
Metric ASHA RASDA Diff.
Train Accuracy $\uparrow$ 0.6179 0.6556 $1.06\times$
Val Accuracy $\uparrow$ 0.6312 0.6348 $1.01\times$
Test Accuracy $\uparrow$ 0.6250 0.6340 $1.01\times$
Runtime (in seconds) $\downarrow$ 17830 11157 $1.60\times$
Table 7: Comparison of RASDA and BOHB on the CV application case, trained for 20 epochs on 64 GPUs on the JURECA-DC-GPU system, evaluating $32$ hyperparameter samples. Results are averaged over three random seeds. Better results ($\uparrow$ or $\downarrow$ depending on the metric) are underlined.
Metric BOHB RASDA Diff.
Train Accuracy $\uparrow$ 0.6130 0.6480 $1.06\times$
Val Accuracy $\uparrow$ 0.6253 0.6271 $1.00\times$
Test Accuracy $\uparrow$ 0.6222 0.6254 $1.01\times$
Runtime (in seconds) $\downarrow$ 16825 11815 $1.42\times$
5.5.2 Comparison to Bayesian Optimization (BO)
Although the primary comparison is between RASDA and its closest scheduling-based counterpart, ASHA, other types of HPO and Neural Architecture Search (NAS) tools are also relevant. One commonly used method is Bayesian Optimization (BO). A comparison between RASDA and the BOHB algorithm on the CV application case is provided below. To function effectively, BO requires the ability to generate new hyperparameter configurations based on past evaluations. As shown in Tab. 7, when both methods are evaluated with the same number of hyperparameter samples, RASDA outperforms BOHB in terms of accuracy and runtime. These results suggest that RASDA achieves a more favorable trade-off between runtime and solution quality compared to traditional BO-based approaches.
@@ 6 Summary and Outlook
In addition, this study represents the first application of systematic HPO to a scientific dataset at the TB scale. A comparison of the application of RASDA on 128 and 1024 GPUs revealed a significant improvement in model performance. Specifically, the larger-scale application identifies a model that is significantly superior in solution quality. These results demonstrate the scalability and efficiency of the RASDA method, thus paving the way for the application of HPO methods on current and future Exascale supercomputers. It should be noted, however, that RASDA relies heavily on efficient distributed training. As such, its full benefits are realized primarily on HPC systems equipped with accelerators such as GPUs. On non-accelerated systems or in settings with very limited communication bandwidth, potential inefficiencies may limit performance.
For future work, the optimal timing for scaling the batch size (through the addition of more GPUs to the data-parallel training loop) should be investigated more thoroughly. Furthermore, the impact of scaling the batch size on various hyperparameters beyond the learning rate (such as the weight decay values) warrants a deeper exploration.